# DCA: Graph-Guided Deep Embedding Clustering for Brain Atlases

### v1

@@ Abstract
Brain atlases are essential for reducing the dimensionality of neuroimaging data and enabling interpretable analysis. However, most existing atlases are predefined, group-level templates with limited flexibility and resolution. We present Deep Cluster Atlas (DCA), a graph-guided deep embedding clustering framework for generating individualized, voxel-wise brain parcellations. DCA combines a pretrained autoencoder with spatially regularized deep clustering to produce functionally coherent and spatially contiguous regions. Our method supports flexible control over resolution and anatomical scope, and generalizes to arbitrary brain structures. We further introduce a standardized benchmarking platform for atlas evaluation, using multiple large-scale fMRI datasets. Across multiple datasets and scales, DCA outperforms state-of-the-art atlases, improving functional homogeneity by 98.8% and silhouette coefficient by 29%, and achieves superior performance in downstream tasks such as autism diagnosis and cognitive decoding. Codes and models will be released soon.
@@ 1 Introduction
Clustering is a cornerstone of brain atlas construction, but off-the-shelf algorithms—such as K-Means Bellec2009MultilevelBA , hierarchical clustering felleman1991distributed , and vanilla spectral methods arslan2015joint —are ill-suited to the characteristics of fMRI data. First, the inherently low signal-to-noise ratio of fMRI hampers these methods’ ability to recover clear boundaries as the number of parcels grows. Second, even a gray-matter mask may contain tens of thousands of voxels, making the computation and storage of a full functional-connectivity matrix prohibitive. Finally, standard clustering optimizes only functional similarity and lacks any notion of spatial continuity, yielding fragmented or anatomically implausible parcels. Although one can incorporate distance-based penalties to encourage spatial contiguity, such strategies demand careful tuning lest they compromise the atlas’s functional coherence.
@@ 2 Related work
Atlas generation based on magnetic resonance imaging (MRI) is the predominant approach for constructing parcellations. Widely adopted algorithms include k-means Bellec2009MultilevelBA , hierarchical clustering felleman1991distributed , spectral clustering arslan2015joint , community detection gordon2017individual , normalized cuts Craddock2012AWB , and statistical learning methodsZhi2023AHB . These techniques aim to maximize within-parcel homogeneity and minimize between-parcel similarity. However, functionally similar voxels are not always spatially contiguous, so enforcing spatial coherence remains challenging. To address this, spatial regularization strategies—such as Markov random field priors schaefer2018local , spatially weighted clustering Craddock2012AWB , and deep Boltzmann machine frameworks Zhi2023AHB —have been introduced to ensure that resulting parcels are both functionally coherent and anatomically contiguous.
@@
Traditional clustering methods measure similarity directly in the original data space, and even when manifold-based techniques are employed, feature extraction and clustering remain two disjointed stagestan2016introduction ; johnson1967hierarchical ; von2007tutorial ; fortunato2016community . Deep clustering, by contrast, integrates these steps. It jointly refines encoder parameters and cluster centers through an auxiliary target distribution derived from the current soft assignments Xie2015UnsupervisedDE ; Cai2022EfficientDE . Beyond its success in image clustering, deep clustering has proven effective for time-series segmentation Lee2024DeepTC , cell detection Abousamra2021MultiClassCD , and disease discovery Yang2024GeneSGANDD . When constructing a brain atlas, both functional similarity and spatial continuity must be preserved. In our framework, each voxel of a single subject serves as an individual sample for deep clustering, and we incorporate a graph-guided spatial prior to ensure contiguous parcels. To our knowledge, this is the first use of deep clustering for generating a fully continuous, voxel-level brain atlas.
@@
To some extent, brain atlas construction shares conceptual similarities with semantic segmentation of brain Shaker2022UNETRDI ; Hatamizadeh2022SwinUS ; Pang2023SlimUS , as both aim to partition the brain into meaningful regions. However, key differences distinguish the two tasks. First, they differ in granularity and objective: brain segmentation typically categorizes voxels based on tissue types—such as whole tumor, tumor core, and enhancing tumor—whereas brain atlases delineate functionally relevant areas, such as precental gyrus, thalamus, hippocampus. Second, semantic segmentation is generally a supervised learning task with well-defined ground truth labels, while atlas construction is inherently unsupervised and must be evaluated using more complex criteria, such as functional or structural homogeneity. Although segmentation techniques have been employed to map existing atlases, they rely heavily on predefined templates and primarily serve to replicate rather than discover novel parcellations. For example, DDparcel assigns voxel-wise labels for 101 anatomical regions based on the Desikan–Killiany atlas, effectively reconstructing rather than redefining an atlas Zhang2023DDParcelDL .
Figure 2: (A) Self-supervised pretraining of Swin-UNETR on group fMRI data: 80% of each volume is Random continuous masked in space and time and reconstructed across 1,000 resting-state trials. Fire icon denotes trainable encoder–decoder weights. (B) Personalized atlas generation at voxel resolution. Individual fMRI volumes are passed through a pretrained encoder to extract both local and global embeddings. Each voxel (red cube) is softly assigned to one of $K$ learnable centroids by measuring its distance to every centroid (top row). Simultaneously, a given ROI mask defines the nodes of a 26-neighborhood graph, whose edge weights are set to the cosine similarity between adjacent embeddings. Sparse spectral clustering on this graph produces hard auxiliary labels (bottom row). A KL-divergence loss then aligns the soft assignments with these auxiliary targets, updating both the centroids and the encoder’s final projection layer. By iterating between refining embeddings, re-weighting graph edges, and updating cluster centers, the method converges to spatially contiguous, functionally coherent parcellations.
@@ 3.2 Personalized atlas generation
We introduce DCA (Deep Clustering Atlas), a self-supervised framework that turns a pretrained 4D encoder into a spatially consistent whole-brain fMRI parcellator (Fig. 2). Firstly, we pretrain a Swin-UNETR autoencoder on masked 4D fMRI blocks (80% spatiotemporal masking), reconstructing the missing voxels. The encoder preserves the full spatial dimensions $(H\times W\times D)$, producing voxel-level embeddings that capture both local and global context (Fig. 2A). The region of interest is defined by any ROI mask. From these embeddings belong to the mask, we maintain $K$ cluster centroids. Each voxel’s embedding is converted into a soft assignment by measuring its distance to every centroid (top row, Fig. 2B). In parallel, we build a 26-nearest-neighbour graph over the ROI masked voxels, weighting edges by pairwise embedding correlations. A graph-cut then yields hard auxiliary targets that enforce spatial contiguity (bottom row, Fig. 2B). Finally, we minimize the KL divergence between the soft assignments and these hard targets, backpropagating through both the encoder’s final layers and the centroids. By alternating between updating embeddings and regenerating auxiliary targets, DCA converges to a voxel-level atlas that is both functionally meaningful and spatially contiguous.
@@
To extract meaningful features from high-dimensional fMRI data, we adopt Swin UNETR as the backbone for pretraininghatamizadeh2021swin (Fig. 2A). Swin UNETR is built upon the Swin Transformer architecture, which introduces a hierarchical structure with shifted windows to compute self-attention in a local and spatially-aware mannerliu2021swin . Unlike standard Vision Transformers (ViT) that operate on flattened global patches and often ignore local continuitydosovitskiy2020image , Swin Transformers preserve spatial hierarchies and are thus better suited for voxel-level modeling in neuroimaging. This property is particularly advantageous for fMRI, where fine-grained spatial relationships between voxels are critical.
We perform self-supervised pretraining using a masked reconstruction objective. During pretraining, we randomly mask out contiguous blocks in both space and time. Specifically, we divide the fMRI volume into non-overlapping spatiotemporal patches and then zero out 80% of those patches along the chosen spatial or temporal axis (more details in Appendix). By forcing the encoder to reconstruct missing segments using information from the surrounding unmasked regions, the model learns representations that capture both local detail and long-range dependencies across space and time. Using Swin UNETR’s encoder-decoder structure in this masked auto-encoding framework, we obtain strong voxel-wise embeddings that serve as the foundation for subsequent clustering and parcellation.
@@
Let $M\in\{0,1\}^{H\times W\times D}$ be a binary mask defining the region of interest, and let $\mathcal{V}=\{i\mid M_{i}=1\}$ index the $N=|\mathcal{V}|$ non-background voxels. We parameterize $K$ cluster centroids as a trainable matrix $\{\boldsymbol{\mu}_{j}\}_{j=1}^{K}\subset\mathbb{R}^{d}$, initialized with orthogonal rows and L2–normalized so that $\boldsymbol{\mu}\,\boldsymbol{\mu}^{\top}=\mathbf{I}$. During each forward pass, we extract the non–background voxel embeddings $\{\mathbf{z}_{i}\}_{i\in V}$ from the Swin-UNETR encoder, compute their Euclidean distances to all centroids, and convert these distances into soft assignments:
@@
The vertex $\mathcal{V}$ is defined by the binary mask $M$. We extract embeddings $\{\mathbf{z}_{i}\in\mathbb{R}^{d}\}_{i\in\mathcal{V}}$ from the pretrained, fine-tunable Swin-UNETR encoder. A 26-neighbourhood graph $G=(V,E)$ , which includes all voxels in a 3×3×3 cube excluding the center, is then constructed by linking each $i\in\mathcal{V}$ to its up to 26 spatial neighbours $j\in\mathcal{V}$, yielding an edge-index array $E\in\mathbb{N}^{2\times|E|}$. Edge weights are given by the cosine similarity of demeaned embeddings:
@@
On the weighted graph $G$, we compute hard auxiliary labels $\mathbf{p}\in\{1,\ldots,K\}^{|\mathcal{V}|}$ via sparse spectral clustering. We form the unnormalized Laplacian L=DWL=D-W, extract the $K$ eigenvectors corresponding to the smallest eigenvalues of $L$, and finally apply K-Means on the resulting $N\times K$ embedding to assign each voxel $i$ its auxiliary label $p_{i}$. To preserve cluster identity across iterations, we realign each new $p_{i}$ to the previous labels via the Hungarian algorithm, yielding an optimal one-to-one mapping. For each iteration, given the previous labels $p_{\text{prev}}$ and the newly obtained labels $p_{\text{new}}$, we first build a cost matrix $C\in\mathbb{Z}^{K\times K}$ where
@@
via the Hungarian algorithm kuhn1955hungarian on C-C, which yields an optimal one-to-one mapping $\pi:\{1,\dots,K\}\to\{1,\dots,K\}$. Finally, we relabel the new clusters $\hat{p}_{i}$ according to $\pi$ by $\hat{p}_{i}=\pi(p_{i})$, thereby preserving label correspondence with earlier iterations.
@@
Let $\mathbf{Q}\in(0,1)^{N\times K}$ be the soft-assignment matrix whose $i$th row is $\mathbf{q}_{i}$, and let $\mathbf{P}\in\{0,1\}^{N\times K}$ be the one-hot encoding of the aligned auxiliary labels $\hat{p}_{i}$. We optimize the Kullback–Leibler divergence
@@
Gradients are back-propagated only to the centroids $\{\boldsymbol{\mu}_{j}\}$ and the final projection block of Swin-UNETR; all other encoder weights remain fixed. This procedure jointly refines both the encoder’s output and the cluster centroids to produce functionally coherent, voxel-level parcellations.
@@
To facilitate downstream use and fair comparison, we also developed a streamlined procedure for deriving a group-level atlas from individual parcellations. We construct the group-level atlas in three steps as detailed in Appendix. First, we pick $K$ template label vectors, each capturing the voxel assignments of one parcel in subjects. Next, we assign every gray-matter voxel to the template vector with which it has highest label similarity. Finally, to guarantee spatial contiguity, we keep only each parcel’s largest connected component and reassign any smaller, isolated regions to adjacent parcels based on local similarity. This yields $K$ contiguous, functionally coherent parcels.
@@ 3.3 Atlas evaluation
To comprehensively assess the quality of any candidate atlas, we provide an interactive evaluation playground. Each input atlas is first spatially normalized to a common template space and resampled to match the reference resolution. We then perform a suite of quantitative tests, including intra-cluster similarity metrics and performance on downstream tasks. This framework enables researchers to compare and validate parcellations across multiple criteria in a standardized environment.
@@ 4 Implementation details
We pretrain our model using a masked reconstruction objective on fMRI data blocks of size $96\times 96\times 96\times 300$, representing 3D spatial volumes with 300 temporal frames. The model is trained for 8 epochs on 2 NVIDIA A100 GPUs using a batch size of 4. The optimizer and learning rate is Adam and 0.01 respectively for both pretraining and fine-tuning. During pretraining, we adopt a masking ratio of 0.8, randomly masking 80% of the input in both spatial and temporal dimensions. The temporal length of the internal representation of Swin-UNETR is downsampled to $T=48$, and the encoder produces a feature map of shape $96\times 96\times 96\times 256$ for clustering, $d$ is 256.

### v2

@@ Abstract
Brain atlases are essential for reducing the dimensionality of neuroimaging data and enabling interpretable analysis. However, most existing atlases are predefined, group-level templates with limited flexibility and resolution. We present Deep Cluster Atlas (DCA), a graph-guided deep embedding clustering framework for generating individualized, voxel-wise brain parcellations. DCA combines a pretrained autoencoder with spatially regularized deep clustering to produce functionally coherent and spatially contiguous regions. Our method supports flexible control over resolution and anatomical scope, and generalizes to arbitrary brain structures. We further introduce a standardized benchmarking platform for atlas evaluation, using multiple large-scale fMRI datasets. Across multiple datasets and scales, DCA outperforms state-of-the-art atlases, improving functional homogeneity by 98.8% and silhouette coefficient by 29%, and achieves superior performance in downstream tasks such as autism diagnosis and cognitive decoding. We also observe that a fine-tuned pretrained model achieves superior results on corresponding task. Codes and models are available at https://github.com/ncclab-sustech/DCA.
@@ 1 Introduction
Clustering is a cornerstone of brain atlas construction, but off‐the‐shelf algorithms—such as K-Means Bellec2009MultilevelBA , hierarchical clustering felleman1991distributed , and vanilla spectral methods arslan2015joint —are ill-suited to the characteristics of fMRI data. First, the inherently low signal-to-noise ratio of fMRI hampers these methods’ ability to recover clear boundaries as the number of parcels grows. Second, even a gray-matter mask may contain tens of thousands of voxels, making the computation and storage of a full functional‐connectivity matrix prohibitive. Finally, standard clustering optimizes only functional similarity and lacks any notion of spatial continuity, yielding fragmented or anatomically implausible parcels. Although one can incorporate distance‐based penalties to encourage spatial contiguity, such strategies demand careful tuning lest they compromise the atlas’s functional coherence.
@@ 2 Related work
Atlas generation based on magnetic resonance imaging (MRI) is the predominant approach for constructing parcellations. Widely adopted algorithms include k‑means Bellec2009MultilevelBA , hierarchical clustering felleman1991distributed , spectral clustering arslan2015joint , community detection gordon2017individual , normalized cuts Craddock2012AWB , and statistical learning methodsZhi2023AHB . These techniques aim to maximize within‑parcel homogeneity and minimize between‑parcel similarity. However, functionally similar voxels are not always spatially contiguous, so enforcing spatial coherence remains challenging. To address this, spatial regularization strategies—such as Markov random field priors schaefer2018local , spatially weighted clustering Craddock2012AWB , and deep Boltzmann machine frameworks Zhi2023AHB —have been introduced to ensure that resulting parcels are both functionally coherent and anatomically contiguous.
@@
Traditional clustering methods measure similarity directly in the original data space, and even when manifold‑based techniques are employed, feature extraction and clustering remain two disjointed stagestan2016introduction ; johnson1967hierarchical ; von2007tutorial ; fortunato2016community . Deep clustering, by contrast, integrates these steps. It jointly refines encoder parameters and cluster centers through an auxiliary target distribution derived from the current soft assignments Xie2015UnsupervisedDE ; Cai2022EfficientDE . Beyond its success in image clustering, deep clustering has proven effective for time‑series segmentation Lee2024DeepTC , cell detection Abousamra2021MultiClassCD , and disease discovery Yang2024GeneSGANDD . When constructing a brain atlas, both functional similarity and spatial continuity must be preserved. In our framework, each voxel of a single subject serves as an individual sample for deep clustering, and we incorporate a graph‐guided spatial prior to ensure contiguous parcels. To our knowledge, this is the first use of deep clustering for generating a fully continuous, voxel‐level brain atlas.
@@
To some extent, brain atlas construction shares conceptual similarities with semantic segmentation of brain Shaker2022UNETRDI ; hatamizadeh2021swin ; Pang2023SlimUS , as both aim to partition the brain into meaningful regions. However, key differences distinguish the two tasks. First, they differ in granularity and objective: brain segmentation typically categorizes voxels based on tissue types—such as whole tumor, tumor core, and enhancing tumor—whereas brain atlases delineate functionally relevant areas, such as precental gyrus, thalamus, hippocampus. Second, semantic segmentation is generally a supervised learning task with well-defined ground truth labels, while atlas construction is inherently unsupervised and must be evaluated using more complex criteria, such as functional or structural homogeneity. Although segmentation techniques have been employed to map existing atlases, they rely heavily on predefined templates and primarily serve to replicate rather than discover novel parcellations. For example, DDparcel assigns voxel-wise labels for 101 anatomical regions based on the Desikan–Killiany atlas, effectively reconstructing rather than redefining an atlas Zhang2023DDParcelDL .
Figure 2: (A) Self-supervised pretraining of Swin-UNETR on group fMRI data: 80% of each volume is Random continuous masked in space and time and reconstructed across 1,000 resting-state trials. Fire icon denotes trainable encoder–decoder weights. (B) Personalized atlas generation at voxel resolution. Individual fMRI volumes are passed through a pretrained encoder to extract both local and global embeddings. Each voxel (red cube) is softly assigned to one of $K$ learnable centroids by measuring its distance to every centroid (top row). Simultaneously, a given ROI mask defines the nodes of a 26‐neighborhood graph, whose edge weights are set to the cosine similarity between adjacent embeddings. Sparse spectral clustering on this graph produces hard auxiliary labels (bottom row). A KL‐divergence loss then aligns the soft assignments with these auxiliary targets, updating both the centroids and the encoder’s final projection layer. By iterating between refining embeddings, re‐weighting graph edges, and updating cluster centers, the method converges to spatially contiguous, functionally coherent parcellations.
@@ 3.2 Personalized atlas generation
We introduce DCA (Deep Clustering Atlas), a self-supervised framework that turns a pretrained 4D encoder into a spatially consistent whole-brain fMRI parcellator (Fig. 2). Firstly, we pretrain a Swin-UNETR autoencoder on masked 4D fMRI blocks (80% spatiotemporal masking), reconstructing the missing voxels. The encoder preserves the full spatial dimensions $(H\times W\times D)$, producing voxel-level embeddings that capture both local and global context (Fig. 2A). The region of interest is defined by any ROI mask. From these embeddings belong to the mask, we maintain $K$ cluster centroids. Each voxel’s embedding is converted into a soft assignment by measuring its distance to every centroid (top row, Fig. 2B). In parallel, we build a 26-nearest-neighbour graph over the ROI masked voxels, weighting edges by pairwise embedding correlations. A graph‐cut then yields hard auxiliary targets that enforce spatial contiguity (bottom row, Fig. 2B). Finally, we minimize the KL divergence between the soft assignments and these hard targets, backpropagating through both the encoder’s final layers and the centroids. By alternating between updating embeddings and regenerating auxiliary targets, DCA converges to a voxel-level atlas that is both functionally meaningful and spatially contiguous.
@@
To extract meaningful features from high-dimensional fMRI data, we adopt Swin-UNETR as the backbone for pretraininghatamizadeh2021swin (Fig. 2A). Swin-UNETR is built upon the Swin Transformer architecture, which introduces a hierarchical structure with shifted windows to compute self-attention in a local and spatially-aware mannerliu2021swin . Unlike standard Vision Transformers (ViT) that operate on flattened global patches and often ignore local continuitydosovitskiy2020image , Swin Transformers preserve spatial hierarchies and are thus better suited for voxel-level modeling in neuroimaging. This property is particularly advantageous for fMRI, where fine-grained spatial relationships between voxels are critical.
We perform self-supervised pretraining using a masked reconstruction objective. During pretraining, we randomly mask out contiguous blocks in both space and time. Specifically, we divide the fMRI volume into non-overlapping spatiotemporal patches and then zero out 80% of those patches along the chosen spatial or temporal axis (more details in Appendix). By forcing the encoder to reconstruct missing segments using information from the surrounding unmasked regions, the model learns representations that capture both local detail and long-range dependencies across space and time. Using Swin-UNETR’s encoder-decoder structure in this masked auto-encoding framework, we obtain strong voxel-wise embeddings that serve as the foundation for subsequent clustering and parcellation.
@@
Let $M\in\{0,1\}^{H\times W\times D}$ be a binary mask defining the region of interest, and let $\mathcal{V}=\{i\mid M_{i}=1\}$ index the $N=|\mathcal{V}|$ non‐background voxels. We parameterize $K$ cluster centroids as a trainable matrix $\{\boldsymbol{\mu}_{j}\}_{j=1}^{K}\subset\mathbb{R}^{d}$, initialized with orthogonal rows and L2–normalized so that $\boldsymbol{\mu}\,\boldsymbol{\mu}^{\top}=\mathbf{I}$. During each forward pass, we extract the non–background voxel embeddings $\{\mathbf{z}_{i}\}_{i\in V}$ from the Swin-UNETR encoder, compute their Euclidean distances to all centroids, and convert these distances into soft assignments:
@@
The vertex $\mathcal{V}$ is defined by the binary mask $M$. We extract embeddings $\{\mathbf{z}_{i}\in\mathbb{R}^{d}\}_{i\in\mathcal{V}}$ from the pretrained, fine‐tunable Swin‐UNETR encoder. A 26‐neighbourhood graph $G=(V,E)$ , which includes all voxels in a 3×3×3 cube excluding the center, is then constructed by linking each $i\in\mathcal{V}$ to its up to 26 spatial neighbours $j\in\mathcal{V}$, yielding an edge‐index array $E\in\mathbb{N}^{2\times|E|}$. Edge weights are given by the cosine similarity of demeaned embeddings:
@@
On the weighted graph $G$, we compute hard auxiliary labels $\mathbf{p}\in\{1,\ldots,K\}^{|\mathcal{V}|}$ via sparse spectral clustering. We form the unnormalized Laplacian L=DWL=D-W, extract the $K$ eigenvectors corresponding to the smallest eigenvalues of $L$, and finally apply K-Means on the resulting $N\times K$ embedding to assign each voxel $i$ its auxiliary label $p_{i}$. To preserve cluster identity across iterations, we realign each new $p_{i}$ to the previous labels via the Hungarian algorithm, yielding an optimal one‐to‐one mapping. For each iteration, given the previous labels $p_{\text{prev}}$ and the newly obtained labels $p_{\text{new}}$, we first build a cost matrix $C\in\mathbb{Z}^{K\times K}$ where
@@
via the Hungarian algorithm kuhn1955hungarian on C-C, which yields an optimal one‐to‐one mapping $\pi:\{1,\dots,K\}\to\{1,\dots,K\}$. Finally, we relabel the new clusters $\hat{p}_{i}$ according to $\pi$ by $\hat{p}_{i}=\pi(p_{i})$, thereby preserving label correspondence with earlier iterations.
@@
Let $\mathbf{Q}\in(0,1)^{N\times K}$ be the soft‐assignment matrix whose $i$th row is $\mathbf{q}_{i}$, and let $\mathbf{P}\in\{0,1\}^{N\times K}$ be the one‐hot encoding of the aligned auxiliary labels $\hat{p}_{i}$. We optimize the Kullback–Leibler divergence
@@
Gradients are back‐propagated only to the centroids $\{\boldsymbol{\mu}_{j}\}$ and the final projection block of Swin‐UNETR; all other encoder weights remain fixed. This procedure jointly refines both the encoder’s output and the cluster centroids to produce functionally coherent, voxel‐level parcellations.
@@
To facilitate downstream use and fair comparison, we also developed a streamlined procedure for deriving a group‐level atlas from individual parcellations. We construct the group‐level atlas in three steps as detailed in Appendix. First, we pick $K$ template label vectors, each capturing the voxel assignments of one parcel in subjects. Next, we assign every gray‐matter voxel to the template vector with which it has the highest label similarity. Finally, to guarantee spatial contiguity, we keep only each parcel’s largest connected component and reassign any smaller, isolated regions to adjacent parcels based on local similarity. This yields $K$ contiguous, functionally coherent parcels.
@@ 3.3 Atlas evaluation
To comprehensively assess the quality of any candidate atlas, we provide an interactive evaluation playground. Each input atlas is first spatially normalized to a common template space and resampled to match the reference resolution. We then perform a suite of quantitative tests, including intra‐cluster similarity metrics and performance on downstream tasks. This framework enables researchers to compare and validate parcellations across multiple criteria in a standardized environment.
@@ 4 Implementation details
We pretrain our model using a masked reconstruction objective on fMRI data blocks of size $96\times 96\times 96\times 300$, representing 3D spatial volumes with 300 temporal frames. The model is trained for 8 epochs on 2 NVIDIA A100 GPUs using a batch size of 4. The optimizer and learning rate are Adam and 0.01 respectively for both pretraining and fine-tuning. During pretraining, we adopt a masking ratio of 0.8, randomly masking 80% of the input in both spatial and temporal dimensions. The temporal length of the internal representation of Swin-UNETR is downsampled to $T=48$, and the encoder produces a feature map of shape $96\times 96\times 96\times 256$ for clustering, $d$ is 256.